OpenStudy (anonymous):

solve for the solution of x Please help will do it in the equation so it looks better

OpenStudy (anonymous):

what's the equation?

OpenStudy (anonymous):

\[\frac{ 2 }{ 1-x}=\frac{ 3 }{1+x }\]

OpenStudy (anonymous):

or there is no solution?

OpenStudy (anonymous):

multiply \[\frac{ 2(x+1)(x-1) }{ 1-x }=\frac{ 3(x+1)(x-1)}{x+1}\], so we'll cancel out (x-1) on top and bottom on the left side of the equation and (x+1) on the right side of the equation. we'll have 2(x+1)=3(x-1), then expand it out, we'll have 2x+2=3x-3, add 3 to both side, 2x+5=3x, then subtract 2x from both side, we'll have 5=1x, so x=5